# Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents

### v1

@@ Appendix A Additional Details on Formulation
A.2 Uncertainty Instantiations and Total Uncertainty Expansion.
@@
Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
@@
Appendix B Alternative Uncertainty Measures
Although we have mainly focused on the above three information-theoretic uncertainty measures throughout the paper, there are plenty of alternatives one can consider depending on the problem setups. In this section, we examine some possible alternative uncertainty measures, Rényi entropy, Tsallis entropy, and informational energy.
@@ Appendix B Alternative Uncertainty Measures
These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
Appendix C Extended Context
@@
One might draw an analogy between the agent UQ and belief tracking in Partially Observable Markov Decision Processes (POMDPs) [kaelbling1998planning]. However, the agent theme challenges the standard setups in traditional POMDPs: (1) there is no explicit belief state, but the agent carries an implicit belief in its memory; (2) actions and observations comprise language and structured strings, spanning an effectively infinite space—far from the small discrete spaces in classical POMDP; and (3) it concerns more on the agent’s uncertainty, whereas POMDP concerns mostly on environment uncertainty.
These connections highlight the distinctive edge of agent UQ, while grounding it in a well-established classical problem set.
Appendix D Note on Implementation
While we set our goal to provide a principled vision with a high-level roadmap for agent UQ, we also discuss some implementation recipes in this section to provide useful guidance for practitioners and researchers.
@@ Appendix D Note on Implementation
To realize the conditional information gate, we first need to classify actions to determine whether each action will contribute to total uncertainty reduction for that task or not. As emphasized in Section 3, considering the interactivity of actions is key to accounting for agent uncertainty dynamics, as well as the evidentiality (or factuality) of the action that is solely considered for the existing approaches so far. Therefore, we may want to implement a compound classifier with both interactivity assessment and evidentiality verification. They can be implemented through a rule/syntactic-based, simple verifier [mu2024rule], an LLM judge with specialized prompting, fine-tuning [liu2025compassverifier], or a hybrid of rule-based and LLM-based [peng2025verif].
Action Category Interactivity Scenario-specific Example
Information-gathering interactive Search flight information ($\tau$-bench Airline); retrieve order status ($\tau$-bench Retail); read message box (ToolSandbox)
Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau$-bench Airline); instruct or ask the user to choose cancellation ($\tau$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
Thinking non-interactive Plan flight schedule ($\tau$-bench Airline); consider returning items ($\tau$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
State-changing tool call (write) non-interactive Update flight booking ($\tau$-bench Airline); return/cancel items ($\tau$-bench Retail); Turn on Wifi (ToolSandbox)
Providing final information to user non-interactive Summarize updated flights ($\tau$-bench Airline); summarize returned items ($\tau$-bench Retail); summarize phonebook update results (ToolSandbox)
Table 2: Example action classes in three scenarios from two benchmarks, $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
@@
In our proposed total uncertainty lower bound Eq. 6, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty. The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or the conformal prediction set size [kumar2023conformal].
However, the observation uncertainty and the initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world. If the system allows multiple trials, one can utilize a sampling-consistency-based approach to estimate the observation uncertainty, likewise the traditional black-box UQ methods. If the system does not allow this, one may try to construct a datastore that contains some relevant data from a similar task and leverage nonparameter estimation methods [ren2023outofdistribution, kotelevskii2022nonparametric]. Another line of methods can be establishing a world model to approximate the environment transition and estimate observation uncertainty from it [chae2025web, zhou2025dinowm, anonymous2026dreamphase].
@@
The proposed information gating mechanism features the information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$. While the estimation of mutual information is notorious for its difficulty, recent advancements on neural estimators [belghazi2018mutual, cheng2020club, mukherjee2020ccmi, molavipour2020conditional, gritsai2025mist] shed light on this problem by enabling estimation in high-dimensional, unstructured data space.
@@
As an analogy to the conditional uncertainty reduction process presented in Section 4.2, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix B. Given that Onicescu’s informational energy brings its correlation coefficient as well, we may be able to implement the exact same style information-gating-based method that has the opposite direction to uncertainty reduction dynamics.
Appendix E Action Categorization
Along with the flight booking scenario in Table 1, we extend our action category analysis to two more scenarios, retail and telecommunication tasks, from $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] in Table 2.
Appendix F Missing Proof
Lemma 2 (Restatement of Lemma 1).
@@ Appendix F Missing Proof
Appendix G Discussion and Future Work

### v2

@@ B.1 Expression of Various Agentic Prompting under Our Formulation
B.2 Uncertainty Instantiations and Total Uncertainty Expansion
@@
Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i})$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
@@
Appendix C Alternative Uncertainty Measures
Although we have mainly focused on the above three information-theoretic uncertainty measures throughout the paper, there are plenty of alternatives one can consider depending on the problem setups. In this section, we examine some possible alternative uncertainty measures, Rényi entropy, Tsallis entropy, and informational energy. These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
@@ Appendix C Alternative Uncertainty Measures
Appendix D Extended Context
@@
One might draw an analogy between the agent UQ and belief tracking in Partially Observable Markov Decision Processes (POMDPs) [kaelbling1998planning]. However, the agent theme challenges the standard setups in traditional POMDPs: (1) there is no explicit belief state, but the agent carries an implicit belief in its memory; (2) actions and observations comprise language and structured strings, spanning an effectively infinite space—far from the small discrete spaces in classical POMDP; and (3) it concerns more on the agent’s uncertainty, whereas POMDP concerns mostly on environment uncertainty. These connections highlight the distinctive edge of agent UQ, while grounding it in a well-established classical problem set.
Appendix E Action-Conditional Uncertainty Dynamics Model
Figure 6: Limitation of existing UQ (a) and our suggestion (b). Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure. We recommend expanding a dimension of interest, interactivity, and moving on to tackle the third challenge (Sec, 4.3) by allowing reducible uncertainty modeling of agents.
Motivation.
From a unified view discussed in Section 3.2, we see that most of the existing UQ methods model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn. One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa. However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward. See Fig. 6 (a), a scenario wherein a user does not fully specify their goal upfront. After searching for options, the agent either reasons itself or interacts with the user to delegate decisions. Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence. Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 2). From our pilot experiments (Figure 9 and Figure 10 provided in Appendix F.2), we observed that an action-independent naive aggregation strategy could not faithfully separate between the success and failure groups on average, resulting in full overall performance reported in Table 2, regardless of the types of uncertainty estimator.
Recommendation: conditional uncertainty reduction process.
We therefore need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, explicitly depending on the agent’s interactivity at each step. While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec. 3. Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
E.1 Implementation Sketch: Conditional Uncertainty Reduction Process
We discuss a possible implementation recipe for this conditional uncertainty reduction process to provide useful guidance for practitioners and researchers.
@@ E.1 Implementation Sketch: Conditional Uncertainty Reduction Process
To realize the conditional information gate, we first need to classify actions to determine whether each action will contribute to total uncertainty reduction for that task or not. As emphasized in Section 3, considering the interactivity of actions is key to accounting for agent uncertainty dynamics, as well as the evidentiality (or factuality) of the action that is solely considered for the existing approaches so far. Therefore, we may want to implement a compound classifier with both interactivity assessment and evidentiality verification. They can be implemented through a rule/syntactic-based, simple verifier [mu2024rule], an LLM judge with specialized prompting, fine-tuning [liu2025compassverifier], or a hybrid of rule-based and LLM-based [peng2025verif]. Below are the example prompts one can adopt in $\tau^{2}$-bench (Retail) for LLM-judge in interactivity and evidentiality classification.
Example System Prompt for Interactivity Classification You are classifying customer service agent actions as either: - "interactive": asking questions, requesting clarification, using tools to collect info - "non_interactive": thinking, state-modifying actions, providing final answers Respond as a JSON array of objects with keys: "action_idx" (int), "label" (str), "justification" (str).
Example System Prompt for Evidentiality Classification You are evaluating whether a customer service agent’s actions are grounded in the conversation context and domain policy. For each assistant action, determine: - Are tool call arguments traceable to user-provided information or prior tool results? - Are text claims about policies consistent with the provided policy document? - Are stated facts (account details, prices, statuses) confirmed by tool results in the conversation? Labels: - "evidential": The action is well-grounded in the conversation and policy. - "non_evidential": The action contains fabricated information, hallucinated tool arguments, or claims contradicting the policy. - "uncertain": Cannot determine grounding with available context. Respond as a JSON array of objects with keys: "action_idx" (int), "label" (str), "justification" (str), "issues" (list of str, empty if none).
@@
In the proposed total uncertainty lower bound Eq. 7, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty. The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or verbalized confidence methods [tian2023just, xiong2024can, yang2024verbalized]. However, as mentioned in Section 4.1, all these existing approaches have their own limitations, and one may have to choose an appropriate estimator depending on the situation.
Meanwhile, observation uncertainty and initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world (See Section 4.2). If the system allows multiple trials, one can utilize a sampling-consistency-based approach to estimate the observation uncertainty, likewise the traditional black-box UQ methods. If the system does not allow this, one may try to construct a datastore that contains some relevant data from a similar task and leverage nonparameter estimation methods [ren2023outofdistribution, kotelevskii2022nonparametric]. Another line of methods can be establishing a world model to approximate the environment transition and estimate observation uncertainty from it [chae2025web, zhou2025dinowm, anonymous2026dreamphase]. Figure 8 provides results on an auxiliary LLM-based approximation that mimics world modeling philosophy.
@@
The proposed information gating mechanism features information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and the initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$. While the estimation of mutual information is notorious for its difficulty, recent advancements on neural estimators [belghazi2018mutual, cheng2020club, mukherjee2020ccmi, molavipour2020conditional, gritsai2025mist] shed light on this problem by enabling estimation in high-dimensional, unstructured data space. Besides, LLM-prompting-based mutual information estimation [robertson2025measure] is also emerging as an attractive line of work, though it has yet to provide a solid theoretical foundation.
@@
As an analogy to the conditional uncertainty reduction process presented before, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix C. Given that Onicescu’s informational energy brings its correlation coefficient as well, we may be able to implement the exact same style information-gating-based method that has the opposite direction to uncertainty reduction dynamics.
E.2 Theoretical Analysis
Lemma 1 (Extrema of Information Gating).
@@ E.2 Theoretical Analysis
E.3 Action Categorization
Action Category Interactivity Example
Table 3: Example action classes in Airline booking scenario in $\tau^{2}$-bench. A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty. See Table 4 for extended analyses.
Action Category Interactivity Scenario-specific Example
Information-gathering interactive Search flight information ($\tau^{2}$-bench Airline); retrieve order status ($\tau^{2}$-bench Retail); read message box (ToolSandbox)
Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau^{2}$-bench Airline); instruct or ask the user to choose cancellation ($\tau^{2}$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
Thinking non-interactive Plan flight schedule ($\tau^{2}$-bench Airline); consider returning items ($\tau^{2}$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
State-changing tool call (write) non-interactive Update flight booking ($\tau^{2}$-bench Airline); return/cancel items ($\tau^{2}$-bench Retail); Turn on Wifi (ToolSandbox)
Providing final information to user non-interactive Summarize updated flights ($\tau^{2}$-bench Airline); summarize returned items ($\tau^{2}$-bench Retail); summarize phonebook update results (ToolSandbox)
Table 4: Example action classes in three scenarios from two benchmarks, $\tau^{2}$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
An actual example categorization of the action interactivity in an airline assistant task is provided in Table 3. Besides, a higher abstraction of action category to two more scenarios, retail in $\tau^{2}$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] is provided in Table 4.
Appendix F Experiment
To ground our statements with empirical results, we conducted a small-scale pilot experiment on a real-world agent benchmark in general assistance tasks to evaluate existing UQ techniques on LLM agents.
F.1 Setup
Dataset and task.
$\tau^{2}$-bench [barres2025tau] (the latest version of $\tau$-bench [yao2024tau]) is a representative agent benchmark adopted in frontier LLM evaluation [gemini_team2026gemini31pro], consisting of three different domains – airline, retail, and telecom, where the total number of sample tasks is 50, 114, and 114, respectively. By following gemini_team2026gemini31pro, we experimented with retail and telecom domains. The tasks in the Retail domain are about helping users cancel or modify pending orders, return or exchange delivered orders, modify user addresses, or provide information. The tasks in the telecom domain are about tackling telecommunication issues in mobile devices, organized into roughly three categories: service issues, mobile data issues, and MMS issues.
LLMs and inference details.
We experiment with two modern LLMs: GPT-4.1 [openai2025gpt41] and Kimi-K2.5 [kimi2026] through Microsoft Azure [^4] https://azure.microsoft.com/en-us AI Foundry API endpoints. Most modern proprietary LLMs do not provide the output log probability in multiple inference engine consistency, and GPT-4.1 was a viable choice for us to adopt here. Meanwhile, small-scale open-source models that can be run in local GPUs did not achieve meaningful performance on $\tau^{2}$-bench due to its task complexity, so we stick with a state-of-the-art large-scale open-source model, Kimi-K2.5, through third party platform. We set the temperature to 0.0 during inference, which is the default value of $\tau^{2}$-bench. Due to the resource constraint, we only run a single trial per task in contrast to the four trials conducted in $\tau^{2}$-bench. User simulator LLM was fixed to Kimi-K2.5 for all tasks, and we only varied agent LLM between GPT-4.1 and Kimi-K2.5.
Implementation of uncertainty estimators.
We adopt three different uncertainty/confidence estimators: negative log-likelihood (NLL), Entropy, and verbalized confidence. The NLL and Entropy are calculated per token and averaged across the entire sequence of actions in the trajectory, whereas the verbalized confidence is calculated per turn and also averaged across the entire trajectory. To be specific,
  • $\hat{U}_{\text{NLL}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{U}_{\text{NLL}}(A_{i}|E_{i-1},O_{i-1})=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}-\log P^{1}_{\pi,\mathcal{T}}(A_{i,j}|E_{i-1},O_{i-1})$
  • $\hat{U}_{\text{Ent}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{U}_{\text{Ent}}(A_{i}|E_{i-1},O_{i-1})=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}\mathbb{E}_{A_{i,j}}[-\log P^{5}_{\pi,\mathcal{T}}(A_{i,j}|E_{i-1},O_{i-1})]$
  • $\hat{C}_{\text{verb}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{C}_{\text{verb}}(A_{i}|E_{i-1},O_{i-1},\texttt{vc-prompt})$
where $L_{i}$ denotes the sequence length of $i$-th turn, $P^{K}$ indicates top-K output token distribution (a chosen token probability if $K=1$), and vc-prompt is a prompt for verbalized confidence we designed as below,
System Prompt for Verbalized Confidence [tian2023just] Provide the probability (0.0 to 1.0) that your current response/action is correct for the given conversation history and context so far. Give ONLY the probability, no other words or explanation.
Evaluation details.
Since $\tau^{2}$-bench does not provide turn-level annotation, we conduct trajectory-level evaluation that compares the trajectory-level uncertainty $U(\mathcal{F}_{\leq T})$ with the final task failure 1r1-r (or success $r$ in the case of verbalized confidence). For this, we measure the area under the receiver operating characteristic curve (AUROC), Spearman rank correlation coefficient $\rho$, and Kendall rank correlation coefficient $\tau$. The average reward (success rate) per model-domain pair is also reported in Table 2.
F.2 Extended Results
This subsection provides full analysis results on the observation uncertainty in Figure 8 and uncertainty evolution in Figure 9 and Figure 10. In Figure 8, we compare three different observation uncertainties over all user messages across trajectories: (1) the GT user simulator uncertainty (computed from the output probability of user simulator LLM), (2) estimates driven from the output probability of agent LLM by feeding user messages along with all previous context, and (3) estimates driven from the output probability of an auxiliary LLM by feeding user messages along with all previous context, coupled with a prompt as below.
System Prompt for Observation Distribution Approximator You are an impartial observer monitoring a customer-service conversation between an AI agent and a human user. Your task is to predict the next observation that the agent will receive: 1. A user message - the customer’s natural-language reply, or 2. A tool result - the deterministic output returned by a backend API after the agent issues a tool call. Given the full conversation history up to this point, generate the most likely continuation as if you were the user or the tool backend. Aim to model the distribution of plausible observations as faithfully as possible: assign high probability to expected responses and low probability to surprising ones.
Figure 8: Comparison of observation uncertainty estimates on $\tau^{2}$-bench. We compare distributions over observation uncertainty estimates, i.e., average NLL and Entropy over all text message tokens from the user. Distributions of the ground truth user simulator LLM and that of the agent-approximated one show remarkable deviation, whereas auxiliary LLM prompting-based estimation shows somewhat close distribution with the ground truth.
We observe remarkable deviation between the GT user simulator’s uncertainties and the estimates from the agent LLM in general. However, using an auxiliary LLM as a world model somehow reduces the gap in many cases. Cost-effective auxiliary model-based approximation can be worth exploring.
Next, Figure 9 and 10 compare uncertainty evolutions over turns between success and failure groups of tasks. To compare varying-length trajectories across different tasks, we normalize the turn index in X-axis into a zero-to-one scale. Results show that existing naive uncertainty aggregation methods can not discern the failure group from the success group to a statistically meaningful degree (the shaded areas in the plots denote a single standard deviation confidence interval), both in the middle of and at the end of the trajectory.
Figure 9: Uncertainty evolution of GPT-4.1 across the trajectory on $\tau^{2}$-bench. We visualize the evolution of action uncertainty/confidence estimates over the normalized turn index, averaged over the entire trajectories from two groups: task success and failure. The shade area denotes one standard deviation.
Figure 10: Uncertainty evolution of Kimi-K2.5 across the trajectory on $\tau^{2}$-bench. We visualize the evolution of action uncertainty/confidence estimates over the normalized turn index, averaged over the entire trajectories from two groups: task success and failure. The shade area denotes one standard deviation.
Appendix G Mini-Survey on Agent Benchmark
We did a small-scale survey on LLM-based agent benchmarks released to the public from February 2023 to February 2026, where we sourced the papers mainly from arxiv.org with non-zero citations, and also referred to ML/NLP conference proceedings. We ended up with a non-exhaustive collection of 44 papers, mainly designed for pure LLM agents, while a few of them also support multimodal scenarios of VLM agents. We confine the scope of the survey to the agent’s problem-solving capability rather than other desired properties such as safety and trustworthiness [zhang2024agent, levy2024st, andriushchenko2025agentharm]. We categorize the papers into three classes depending on the evaluation granularity: turn-level, milestone-level, and trajectory-level. Table 5 summarizes the turn-level and milestone-level evaluation benchmarks, and Table 6 summarizes the trajectory-level ones. The scarcity of fine-grained benchmarks is severe and becomes a real bottleneck to establishing and evaluating the agent UQ framework.
Title Citation Evaluation
Agent-as-a-Judge: Evaluate Agents with Agents pmlr-v267-zhuge25a Turn-level
API-Bank: A Comprehensive Benchmark for Tool-Augmented LLMs li-etal-2023-api Turn-level
Clembench: Using Game Play to Evaluate Chat-Optimized Language Models as Conversational Agents chalamalasetti2023clembench Turn-level
Mind2Web: Towards a Generalist Agent for the Web deng2023mind2web Turn-level
AgentBoard: An Analytical Evaluation Board of Multi-turn LLM Agents chang2024agentboard Milestone-level
DevBench: A Comprehensive Benchmark for Software Development li2025prompting Milestone-level
Gaia2: Benchmarking LLM Agents on Dynamic and Asynchronous Environments froger2026gaia2 Milestone-level
PaperBench: Evaluating AI’s Ability to Replicate AI Research pmlr-v267-starace25a Milestone-level
ScienceAgentBench: Toward Rigorous Assessment of Language Agents for Data-Driven Scientific Discovery chen2025scienceagentbench Milestone-level
ST-WebAgentBench: A Benchmark for Evaluating Safety and Trustworthiness in Web Agents levy2024st Milestone-level
The Berkeley Function Calling Leaderboard (BFCL): From Tool Use to Agentic Evaluation of Large Language Models patil2025berkeley Milestone-level
TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks xu2025theagentcompany Milestone-level
ToolSandBox: A Stateful, Conversational, Interactive Evaluation Benchmark for LLM Tool Use Capabilities lu2025toolsandbox Milestone-level
WebCanvas: Benchmarking Web Agents in Online Environments pan2024webcanvas Milestone-level
Table 5: Agent benchmark research with turn-level and milestone-level evaluation protocols.
Title Citation
AgentBench: Evaluating LLMs as Agents liu2024agentbench
AgentGym: Evolving Large Language Model-based Agents across Diverse Environments xi-etal-2025-agentgym
AppWorld: A Controllable World of Apps and People for Benchmarking Interactive Coding Agents trivedi-etal-2024-appworld
AssistantBench: Can Web Agents Solve Realistic and Time-Consuming Tasks? yoran2024assistantbench
BrowseComp: A Simple Yet Challenging Benchmark for Browsing Agents wei2025browsecomp
CORE-Bench: Fostering the Credibility of Published Research Through a Computational Reproducibility Agent Benchmark siegel2024corebench
Finance Agent Benchmark: Benchmarking LLMs on Real-world Financial Research Tasks bigeard2025finance
GAIA: A Benchmark for General AI Assistants mialon2023gaia
GameBench: Evaluating Strategic Reasoning Abilities of LLM Agents costarelli2024gamebench
InterCode: Standardizing and Benchmarking Interactive Coding with Execution Feedback yang2023intercode
MCPAgentBench: A Real-world Task Benchmark for Evaluating LLM Agent MCP Tool Use liu2025mcpagentbench
MCP-Atlas: A Large-Scale Benchmark for Tool-Use Competency with Real MCP Servers bandi2026mcpatlas
MedAgentBench: A Realistic Virtual EHR Environment to Benchmark Medical LLM Agents jiang2025medagentbench
MINT: Evaluating LLMs in Multi-turn Interaction with Tools and Language Feedback wang2024mint
MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation pmlr-v235-huang24y
OSWorld: Benchmarking Multimodal Agents for Open-Ended Tasks in Real Computer Environments xie2024osworld
SafeAgentBench: A Benchmark for Safe Task Planning of Embodied LLM Agents yin2024safeagentbench
SmartPlay: A Benchmark for LLMs as Intelligent Agents wu2024smartplay
SOTOPIA: Interactive Evaluation for Social Intelligence in Language Agents zhou2024sotopia
SWE-bench: Can Language Models Resolve Real-World GitHub Issues? jimenez2024swebench
Terminal-Bench: Benchmarking Agents on Hard, Realistic Tasks in Command Line Interfaces merrill2026terminal
The Tool Decathlon: Benchmarking Language Agents for Diverse, Realistic, and Long-Horizon Task Execution li2025tooldecathlon
ToolLLM: Facilitating Large Language Models to Master 16000+ Real-world APIs qin2024toolllm
TravelPlanner: A Benchmark for Real-World Planning with Language Agents xie2024travelplanner
Vending-Bench: A Benchmark for Long-Term Coherence of Autonomous Agents backlund2025vending
WebArena: A Realistic Web Environment for Building Autonomous Agents zhou2024webarena
WebShop: Towards Scalable Real-World Web Interaction with Grounded Language Agents yao2022webshop
WorkArena: How Capable Are Web Agents at Solving Common Knowledge Work Tasks? drouin2024workarena
$\tau$-bench: A Benchmark for Tool-Agent-User Interaction in Real-World Domains yao2024tau
$\tau^{2}$-Bench: Evaluating Conversational Agents in a Dual-Control Environment barres2025tau
Table 6: Agent benchmark research with a trajectory-level evaluation protocol.
Appendix H Discussion and Future Work